OpenStudy (anonymous):

Can someone help me find the angle between these two vectors? it has to be to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

jimthompson5910 (jim_thompson5910):

use the formula \[\Large \theta = \arccos\left(\frac{u \cdot v}{|u||v|}\right) \]